OpenStudy (anonymous):

(a) In an arithmetic progression, the sum, Sn , of the first n terms is given by Sn=2n^2 + 8n. Find the first term and the common difference of the progression. ans ; a =64 d= 3/4

jhonyy9 (jhonyy9):

so for the first term n=1 what mean that Sn = 10

Parth (parthkohli):

\[a_n = S_n - S_{n - 1}= 2n^2 + 8n - \left(2(n-1)^2 + 8(n-1)\right)\]

Parth (parthkohli):

You know what... I don't think you even need to derive the formula for the \(n\)th term. But you can apply a similar method to find the common difference.\[S_1 = a = 2 + 8 = 10.\]\[S_2 = a + (a + d) = \color{#c00}{2a + d} = 2(2)^2 + 8(2) = \color{#c00}{24}\]The first equation tells us that \(a = 10\), so we'll plug that into\[2a + d = 24\Rightarrow 20 + d = 24 \Rightarrow d = 4\]

OpenStudy (anonymous):

Assume n=1,2,3.. then s1=2.1^2+8.1=10 s2=2.2^2+8.2=24 and s3=2.3^2+8.3=42 1st term=10 2nd term=s2-s1=24-10=14 3rd term=s3-s2=42-24=18 a=10 and d=4 Check, Sn=n/2{2a+(n-1)d} Sn=n/2{2.10+(n-1)4} Sn=n/2(20+4n-4) Sn=n(10+2n-2) Sn=2n^2+8n Done!!!